### v1

@@ Abstract
In the present paper local derivations and local automorphisms of five-dimensional naturally graded nilpotent associative algebras are studied. Namely, a general form of the matrices of local derivations and local automorphisms of algebras $\pi_{2}$ and $\pi_{3}$ is clarified. It turns out that the general form of the matrix of an automorphism (derivation) on these algebras does not coincide with the local automorphism’s (resp. local derivation’s) matrix’s general form on these algebras. Therefore, these associative algebras have local automorphisms (resp. local derivations) that are not automorphisms (resp. derivations). We also establish a relationship between local automorphisms and local derivations via an exponential expression. We prove that the sets of local derivations of algebras $\pi_{2}$ and $\pi_{3}$ form Lie algebras with respect to the Lie brackets. Thus, we show that the Lie algebra problem from the Ayupov-Eldique-Kudaybergenov problems for local derivations of the algebras under consideration has a positive solution. The other problems from the Ayupov-Eldique-Kudaybergenov problems have a positive (negative) solution for algebra $\pi_{2}$ (resp. algebra $\pi_{3}$).
@@ 1 Introduction
In the present paper we show that in the cases of the algebras under consideration the Lie algebra problem (2) from Problems 1 for local derivations has a positive solution. The other problems (1) and (3) from Problems 1 have a positive solution for algebra $\pi_{2}$ and a negative solution for algebra $\pi_{3}$.
@@ 2 Preliminaries
We assume throughout this section that $\mathbb{F}$} is the field of complex numbers. Let $\mathcal{A}$ be a naturally graded $n$-dimensional quasi-filiform associative algebra. Then, there are two possibilities for the characteristic sequence, either $C(\mathcal{A})={(n-2,1,1)}$ or $C(\mathcal{A})=(n-2,2)$}.
@@ 7 The relationship between local derivations and local automorphisms of $\pi_{2}$ and $\pi_{2}$
Now we compute the components of the matrix $B=\exp(A)=E+\sum_{n=1}^{\infty}\frac{A^{n}}{n!}$, where $E$ is the unit matrix. Let $B=(b_{i,j})_{i,j=1}^{5}$. Since, for $x\in{\mathbb{F}}$,
@@
$\frac{(2^{0}+2^{1}+2^{2}+2^{3}+2^{4})x_{1}^{4}}{5!}+\dots+\frac{(\sum_{k=0}^{n}2^{k})x_{1}^{n}}{(n+1)!}+\dots,$
@@
$\frac{((1+2+2^{2}+...+2^{n-2})+3((1+2+2^{2}+...+2^{n-3})+\dots+3((1+2)+3)))x_{1}^{n-2}}{n!}+\dots,$
$b_{3,2}=\lambda_{3,2}(x_{1})x_{5}$, where
@@
$\frac{(2^{n-1}+3_{1}(2^{n-2}+3_{2}(2^{n-3}+3_{3}(+\dots+3_{n-1}))))x_{1}^{n-1}}{n!}+\dots,$
$b_{3,4}=\lambda_{3,4}(x_{1})x_{2}$, where
@@
Similar to the case of algebra $\pi_{3}$, since the group of all matrices of local automorphisms corresponding to the group $LocAut(\pi_{2})$ is not topologically closed, we can not apply Cartan’s theorem to group $LocAut(\pi_{2})$.
Note that:
  • The remaining elements are expressed through $b_{11}$ or are equal to 0.
  • The sign in $\pm b_{11}^{3}$ gives two possible options.
$\mathcal{M}(LocAut(\pi_{3}))=\mathcal{M}(LocAut(\pi_{3}))_{+}\cup\mathcal{M}(LocAut(\pi_{3}))_{-},$
where $\mathcal{M}(LocAut(\pi_{3}))_{+}$ is a subset of matrices with $b_{11}^{3}$, $\mathcal{M}(LocAut(\pi_{3}))_{-}$ is a subset of matrices with $b_{11}^{3}$.
Consider the mapping
$\Phi_{\pm}:\mathbb{R}^{7}\supset U\to\mathbb{R}^{5\times 5},$
which, for each set of parameters
$(b_{11},b_{21},b_{31},b_{32},b_{34},b_{51},b_{54})\mapsto B,$
forms a matrix of the form (15) with a fixed sign in $\pm b_{11}^{3}$. This mapping is smooth (all components are polynomials in the parameters), and its image is a smooth submanifold of dimension $7$ in $\mathbb{R}^{25}$, since the mapping is an embedding. Thus, each of the sets $\mathcal{M}(LocAut(\pi_{3}))_{+}$ and $\mathcal{M}(LocAut(\pi_{3}))_{-}$ is a smooth $7$-dimensional manifold in $\mathbb{R}^{25}$.
The set $\mathcal{M}(LocAut(\pi_{3}))=\mathcal{M}(LocAut(\pi_{3}))_{+}\cup\mathcal{M}(LocAut(\pi_{3}))_{-}$ is the union of two disjoint smooth manifolds. Note that they are disjoint because if $b_{11}\neq 0$, then $b_{11}^{3}\neq-b_{11}^{3}$, which means $\pm b_{11}^{3}$ are distinct. Hence, matrices cannot simultaneously lie in both sets.
The union of two disjoint smooth submanifolds of the same dimension is not a smooth manifold, since it is not connected, and at the ”joint” (which is absent) there is no smooth structure uniting both parts into a single manifold. Therefore, the set of such matrices is not a smooth manifold. Thus, the following theorem holds true.
Theorem 15.
The group $LocAut(\pi_{3})$ of all local automorphisms of algebra $\pi_{3}$ is not a Lie group.
@@
Since a smooth mapping $\phi:\mathbb{R}^{11}\to\mathbb{R}^{25}$ is an embedding (an injective mapping with regular differential), we have the image $\mathcal{M}(LocAut(\pi_{2}))$ is a smooth submanifold of dimension $11$. Thus, the set $\mathcal{M}(LocAut(\pi_{2}))$ is a smooth embedded submanifold of $\mathbb{R}^{25}$ (the space of all $5\times 5$ matrices), of dimension $11$. As the result we have the following theorem.
@@ 8 Conclusion
The vector space $LocDer(\pi_{2})$ ($LocDer(\pi_{3})$) of all local derivations of the algebra $\pi_{2}$ (resp. $\pi_{3}$) forms a Lie algebra with respect to the Lie multiplication $\nabla,\Delta]=\nabla\Delta-\Delta\nabla$.
For every local derivation $\nabla\in LocDer(\pi_{2})$ the mapping $\exp(\nabla)$ is a local automorphism of $\pi_{2}$. For every local derivation $\nabla\in LocDer(\pi_{3})$ the mapping $\exp(\nabla)$ is a local automorphism of $\pi_{3}$ and, conversely, for every local automorphism $\Phi\in LocAut(\pi_{3})$ there exists a local derivation $\nabla\in LocDer(\pi_{3})$ such that $\Phi=\exp(\nabla)$. The group $LocAut(\pi_{2})$ of local automorphisms of algebra $\pi_{2}$ is a Lie group while the group $LocAut(\pi_{3})$ of local automorphisms of algebra $\pi_{3}$ is not.

### v2

@@ Abstract
In the present paper local derivations and local automorphisms of five-dimensional naturally graded nilpotent associative algebras are studied. Namely, a general form of the matrices of local derivations and local automorphisms of algebras $\pi_{2}$ and $\pi_{3}$ is clarified. It turns out that the general form of the matrix of an automorphism (derivation) on these algebras does not coincide with the local automorphism’s (resp. local derivation’s) matrix’s general form on these algebras. Therefore, these associative algebras have local automorphisms (resp. local derivations) that are not automorphisms (resp. derivations). We also establish a relationship between local automorphisms and local derivations via an exponential expression. We prove that the sets of local derivations of algebras $\pi_{2}$ and $\pi_{3}$ form Lie algebras with respect to the Lie brackets. Thus, we show that the Lie algebra problem from the Ayupov-Eldique-Kudaybergenov problems for local derivations of the algebras under consideration has a positive solution. The remaining problems from the Ayupov-Eldique-Kudaybergenov problems also have a positive solution for algebras $\pi_{2}$ and $\pi_{3}$.
@@ 1 Introduction
In the present paper we show that in the cases of the algebras under consideration Problems 1 also have a positive solution.
@@ 2 Preliminaries
We assume throughout this paper that $\mathbb{F}$} is the field $\mathbb{C}$} of complex numbers. Let $\mathcal{A}$ be a naturally graded $n$-dimensional quasi-filiform associative algebra. Then, there are two possibilities for the characteristic sequence, either $C(\mathcal{A})={(n-2,1,1)}$ or $C(\mathcal{A})=(n-2,2)$}.
@@ 7 The relationship between local derivations and local automorphisms of $\pi_{2}$ and $\pi_{2}$
Now we compute the components of the matrix $B=\exp(A)=E+\sum_{n=1}^{\infty}\frac{A^{n}}{n!}$, where $E$ is the unit matrix. Let $B=(b_{i,j})_{i,j=1}^{5}$. Since, for $x\in{\mathbb{C}}$, the infinite sum to the right of the following equality is limited
@@
$\frac{(2^{0}+2^{1}+2^{2}+2^{3}+2^{4})x_{1}^{4}}{5!}+\dots+\frac{(\sum_{k=0}^{n}2^{k})x_{1}^{n}}{(n+1)!}+\dots$
$\sum_{n=0}^{\infty}\frac{(\sum_{k=0}^{n}2^{k})x_{1}^{n}}{(n+1)!}.$
The infinite sum is bounded and converges for all complex values of $x$. Indeed, we simplify the inner sum
$\sum_{k=0}^{n}2^{k}=2^{n+1}-1.$
Now, substitute this result into the original power series
$\lambda_{2,1}(x_{1})=\sum_{n=0}^{\infty}\frac{(2^{n+1}-1)x^{n}}{(n+1)!}.$
We can split this into two separate series
$\lambda_{2,1}(x_{1})=\sum_{n=0}^{\infty}\frac{2^{n+1}x^{n}}{(n+1)!}-\sum_{n=0}^{\infty}\frac{x^{n}}{(n+1)!}.$
The sums we rewrite as follows
$\lambda_{2,1}(x_{1})=\frac{1}{x}\sum_{m=1}^{\infty}\frac{(2x)^{m}}{m!}-\frac{1}{x}\sum_{m=1}^{\infty}\frac{x^{m}}{m!}.$
Then we conclude that
$\lambda_{2,1}(x_{1})=\frac{e^{2x}-1}{x}-\frac{e^{x}-1}{x}=\frac{e^{2x}-e^{x}}{x}.$
Therefore $\lambda_{2,1}(x_{1})\in{\mathbb{C}}$.
@@
$\frac{((\sum_{k=0}^{n-2}2^{k})+3((\sum_{k=0}^{n-3}2^{k})+3((\sum_{k=0}^{n-4}2^{k})+\dots+3((1+2)+3)\dots)))x_{1}^{n-2}}{n!}+\dots.$
The complex infinite sum $\lambda_{3,1}(x)$ is also bounded and converges for all complex values of $x$. Indeed, let
  • $a_{2}=1$
  • $a_{3}=(1+2)+3=6$
  • $a_{4}=(1+2+2^{2})+3(a_{3})=7+18=25$
  • $a_{5}=(1+2+2^{2}+2^{3})+3(a_{4})=15+75=90$ and so on
We can derive the following recursive relation
$a_{n}=(2^{n-1}-1)+3a_{n-1}.$
By solving this linear recurrence, we find the following closed-form expression for $a_{n}$
$a_{n}=\frac{3^{n}-2^{n+1}+1}{2}.$
Substituting $a_{n}$ back into the original sum we have
$\lambda_{3,1}(x)=\sum_{n=2}^{\infty}\frac{a_{n}x^{n-2}}{n!}=\sum_{n=2}^{\infty}\frac{(3^{n}-2\cdot 2^{n}+1)x^{n-2}}{2\cdot n!}.$
Using the Taylor series $e^{x}=\sum_{n=0}^{\infty}\frac{x^{n}}{n!}$, we obtain
$\lambda_{3,1}(x)=\frac{e^{3x}-2e^{2x}+e^{x}}{2x^{2}}=\frac{e^{x}(e^{x}-1)^{2}}{2x^{2}}.$
Therefore, the sum is bounded for all $x\in\mathbb{C}$.
Now we consider $\mu_{3,1}(x_{1})$. We can write the given infinite series in a general summation form:
$\mu_{3,1}(x_{1})=\sum_{n=2}^{\infty}\frac{A_{n}}{n!}x_{1}^{n-2},$
where $A_{2}=1$, $A_{3}=(1+2)+3=6$, $A_{4}=(1+2+4)+3\times 6=25$, $A_{5}=(1+2+4+8)+3\times 25=90$ and so on. From this pattern, we can see that
$A_{n}=3A_{n-1}+(2^{n-1}-1).$
By solving this recurrence equation with the initial condition $A_{2}=1$, we obtain that
$A_{n}=\frac{1}{2}\cdot 3^{n}-2^{n}+\frac{1}{2}.$
Hence,
$\mu_{3,1}(x_{1})=\frac{1}{x_{1}^{2}}\sum_{n=2}^{\infty}\frac{\frac{1}{2}\cdot 3^{n}-2^{n}+\frac{1}{2}}{n!}x_{1}^{n}.$
From this it follows that
$\mu_{3,1}(x_{1})=\frac{e^{x_{1}}(e^{x_{1}}-1)^{2}}{2x_{1}^{2}}.$
Similarly we have
$\lambda_{3,2}(x_{1})=\frac{1}{x_{1}}\sum_{n=1}^{\infty}\frac{(3x_{1})^{n}}{n!}-\frac{1}{x_{1}}\sum_{n=1}^{\infty}\frac{(2x_{1})^{n}}{n!}$
$\frac{e^{3x_{1}}-1}{x_{1}}-\frac{e^{2x_{1}}-1}{x_{1}}=\frac{e^{3x_{1}}-e^{2x_{1}}}{x_{1}},$
where
@@
$\frac{(2^{n-1}+3_{1}(2^{n-2}+3_{2}(2^{n-3}+3_{3}(+\dots+3_{n-1}))))x_{1}^{n-1}}{n!}+\dots$
and $b_{3,2}=\lambda_{3,2}(x_{1})x_{5}$. Therefore, $b_{3,2}\in\mathbb{C}$.
Also we have $b_{3,4}=\lambda_{3,4}(x_{1})x_{2}=\frac{e^{3x_{1}}-e^{x_{1}}}{2x_{1}}x_{2}\in\mathbb{C}$, where
@@
Since the groups of matrices of local automorphisms $LocAut(\pi_{2})$ and $LocAut(\pi_{3})$ are not topologically closed, we can not apply Cartan’s theorem to the groups $LocAut(\pi_{2})$, $LocAut(\pi_{3})$.
@@
Since a smooth mapping $\phi:\mathbb{R}^{11}\to\mathbb{R}^{25}$ is an embedding (an injective mapping with regular differential), we have the image $\mathcal{M}(LocAut(\pi_{2}))$ is a smooth submanifold of dimension $11$. Thus, the set $\mathcal{M}(LocAut(\pi_{2}))$ is a smooth embedded submanifold of $\mathbb{R}^{25}$ (the space of all $5\times 5$ matrices), of dimension $11$. So $LocAut(\pi_{2})$ is a Lie group and the tangent space $Lie(LocAut(\pi_{2}))$ coincides with $LocDer(\pi_{2})$. As the result we have the following theorem.
@@
Note that:
  • The remaining elements are expressed through $b_{11}$ or are equal to 0.
  • The sign in $\pm b_{11}^{3}$ gives two possible options.
$\mathcal{M}(LocAut(\pi_{3}))=\mathcal{M}(LocAut_{+}(\pi_{3}))\cup\mathcal{M}(LocAut_{-}(\pi_{3})),$
where $\mathcal{M}(LocAut_{+}(\pi_{3}))$ is a subset of matrices with $b_{11}^{3}$, $\mathcal{M}(LocAut_{-}(\pi_{3}))$ is a subset of matrices with $b_{11}^{3}$.
Consider the mapping
$\Phi_{\pm}:\mathbb{R}^{7}\supset U\to\mathbb{R}^{5\times 5},$
which, for each set of parameters
$(b_{11},b_{21},b_{31},b_{32},b_{34},b_{51},b_{54})\mapsto B,$
forms a matrix of the form (15) with a fixed sign in $\pm b_{11}^{3}$. This mapping is smooth (all components are polynomials in the parameters), and its image is a smooth submanifold of dimension $7$ in $\mathbb{R}^{25}$, since the mapping is an embedding. Thus, each of the sets $\mathcal{M}(LocAut_{+}(\pi_{3}))$ and $\mathcal{M}(LocAut_{-}(\pi_{3}))$ is a smooth $7$-dimensional manifold in $\mathbb{R}^{25}$.
The set $\mathcal{M}(LocAut(\pi_{3}))=\mathcal{M}(LocAut_{+}(\pi_{3}))\cup\mathcal{M}(LocAut_{-}(\pi_{3}))$ is the union of two disjoint smooth manifolds. Note that they are disjoint because if $b_{11}\neq 0$, then $b_{11}^{3}\neq-b_{11}^{3}$, which means $\pm b_{11}^{3}$ are distinct. Hence, matrices cannot simultaneously lie in both sets.
The union of two disjoint smooth submanifolds is a smooth manifold, Therefore, the set of such matrices is a smooth manifold. So $LocAut(\pi_{3})$ is a Lie group. Thus, the following theorem holds true.
Theorem 16.
The group $LocAut(\pi_{3})$ of all local automorphisms of algebra $\pi_{3}$ is a Lie group.
@@ 8 Conclusion
The vector space $LocDer(\pi_{2})$ ($LocDer(\pi_{3})$) of all local derivations of algebra $\pi_{2}$ (resp. $\pi_{3}$) forms a Lie algebra with respect to the Lie multiplication $\nabla,\Delta]=\nabla\Delta-\Delta\nabla$.
For every local derivation $\nabla\in LocDer(\pi_{2})$ the mapping $\exp(\nabla)$ is a local automorphism of $\pi_{2}$. For every local derivation $\nabla\in LocDer(\pi_{3})$ the mapping $\exp(\nabla)$ is a local automorphism from $LocAut_{+}(\pi_{3})$ and, conversely, for every local automorphism $\Phi\in LocAut_{+}(\pi_{3})$ there exists a local derivation $\nabla\in LocDer(\pi_{3})$ such that $\Phi=\exp(\nabla)$. The groups $LocAut(\pi_{2})$ and $LocAut(\pi_{3})$ of local automorphisms of algebras $\pi_{2}$ and $\pi_{3}$ respectively are Lie groups.